# Changing the ranking in eigenvector centrality of a weighted graph by small perturbations

### v1

@@ 5.2. Constrained gradient of the functional
with an appropriate matrix-valued function $Z_{\varepsilon}$, which - in a natural way - is interpreted as the constrained gradient of $F_{\varepsilon}$ (with respect to the Frobenius inner product). Projecting it onto the space of real matrices with sparsity pattern ${\mathcal{E}}$ (first constrain) and unit norm (second constrain), we obtain the sought fully constrained gradient.
@@
Lemma 5.1 (Fully constrained gradient of the functional).
@@
(13) $\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=\eta\Big{\langle}\widehat{G}_{\varepsilon}(E),\dot{E}\Big{\rangle},\qquad\widehat{G}_{\varepsilon}(E)=G_{\varepsilon}(E)-\big{\langle}G_{\varepsilon}(E),E\big{\rangle}E,$
where $\eta>0$ is a suitable scaling factor and $\widehat{G}_{\varepsilon}(E)$ is the fully constrained gradient of $F_{\varepsilon}$ with
(14) $G_{\varepsilon}(E)=P_{\mathcal{E}}\left(\sum_{k=1}^{m}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\Bigl{(}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)(A_{\varepsilon}^{\#})^{\top}\mathrm{v}\mathrm{v}^{\top}-(A_{\varepsilon}^{\#})^{\top}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}\Bigr{)}\right)$
the structure constrained gradient, having the sparsity pattern determined by the set of edges $\mathcal{E}$.
@@
Lemma 4.1 applied to $A+\varepsilon E(t)$ yields
@@
This gives the structure-constrained gradient $G_{\varepsilon}(E)=P_{\mathcal{E}}(R)$ (see (14)).
In order to include the further norm constraint (ia) we project the gradient onto the tangent space to the set $\mathcal{S}_{1}=\{M\in\mathbb{R}^{n\times n}\ |\ \|M\|=1\}$ at $E$. To comply with the constraint ${\|E(t)\|^{2}=1}$, we must have
(16) $0=\frac{1}{2}\,\frac{d}{dt}\|E(t)\|^{2}=\langle E(t),\dot{E}(t)\rangle.$
We are thus led to the constrained optimization problem $\arg\min\limits_{\langle E,Z\rangle=0}\ \langle G,Z\rangle$ (where $G=G_{\varepsilon}(E)$). In order to obtain a unique solution we also impose that the solution (which represents a direction) has unit norm. Thus we have to solve the problem
(17) $Z_{\star}$ $=$ $\arg\min_{\|Z\|=1,\,\langle E,Z\rangle=0}\ \langle G,Z\rangle.$
Its solution is given by
(18) $\mu Z_{\star}$ $=G-\langle E,G\rangle\,E,$
where $\mu$ is the Frobenius norm of the matrix on the right-hand side. The solution is unique if $G$ is not a multiple of $E$. The result follows on noting that the real inner product with a given vector (which here is a matrix) is maximized over a subspace by orthogonally projecting the vector onto that subspace. The expression in (18) is the orthogonal projection of $G$ onto the orthogonal complement of the span of $E$, which is the tangent space at $E$ of the manifold of matrices of unit Frobenius norm. Setting $\widehat{G}_{\varepsilon}(E)$ as the right-hand-side of (18) gives (13). This concludes the proof. ∎
Constrained gradient system
In order to minimize the functional $F_{\varepsilon}$ over the space of $\varepsilon$-feasible matrices it is natural to consider the associated gradient system
@@
(19) $\dot{E}=-\widehat{G}_{\varepsilon}(E)=-G_{\varepsilon}(E)+\gamma E\quad\hbox{ with }\quad\gamma=\langle E,G_{\varepsilon}(E)\rangle.$
It may happen that along the solution trajectory of (19) some positive entry of $A+\varepsilon E$ may reach the small positive value $\delta$. In our experiments we never observed such an occurrence, although this is a potential situation. In subsection 5.4 we comment on how to deal with this further constraint.
Thus we treat the constraint (ii) (which is a relaxation of non-negativity) by penalization. Omitting for now penalization (which we shall discuss later), we analyze in the next section the peculiar properties of the system of ODEs (19).
@@ 5.3. Monotonicity and stationary points
The following monotonicity result follows directly from the construction of the gradient system (19) (for the moment we assume that the non-negativity constraint does not activate so that penalization is not necessary, which is what we observe in all the tests we have done).
@@
Furthermore, the following result characterizes stationary points.
Theorem 5.2.
The following statements are equivalent along solutions of (19):
@@
    $E$ is a real multiple of $G_{\varepsilon}(E)$.
@@
Using Lemma 5.1 and (19) we obtain, with $G=G_{\varepsilon}(E)$,
(20) $\frac{1}{\varepsilon}\,\dot{F}_{\varepsilon}=\langle G,\dot{E}\rangle=\langle G,-G+\gamma E\rangle=-\|G\|^{2}+\langle G,E\rangle^{2}\leq 0.$
By (20), we have that
$\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=0\qquad\Longrightarrow\qquad\mbox{$E$ is a multiple of $G$}.$
Then $\langle G,\dot{E}\rangle$ with $\dot{E}=-G+\gamma E$ implies $\dot{E}=0$.
Clearly, 3. implies 2., which implies 1. Since the Cauchy–Schwarz inequality in (20) is strict unless $E$ is a multiple of $G$, we obtain that 1. implies 3. ∎
@@ 5.4. Underlying low-rank property of extremizers
For a sparse graph we have a number of ODEs which is given by the nonzero entries of the sparsity pattern associated to the graph. However, for a dense graph we may have up to $n^{2}$ ODEs. Following the lines of [19], we are able to exploit the underlying low-rank structure of the solutions of the optimization problem we consider. By Theorem 5.1 we have that the stationary points which we aim to compute are proportional to the gradient $G$, which - as we have seen - is the projection of a rank-$1$ matrix. Proceeding similarly to [19], by writing $E=P_{\mathcal{E}}(H)$ for a rank-$1$ matrix $H$, we are able to obtain a new system of ODEs for $H$, whose dynamics is constrained to the manifold of rank-$1$ matrices. We can prove that the stationary points of the two systems of ODEs are in a one-to-one correspondence and also that the system of ODEs for $H$, even if it loses the gradient system structure, converges - at least locally - to the stationary points. We omit a full derivation for sake of conciseness and refer the reader to [19] for details.
5.5. Non-negativity constraint
A further constraint on $E(t)$ is relaxed non-negativity, which means $(A+\varepsilon E(t))_{ij}\geq\delta$ (with $\delta$ a very small tolerance) for all $t$.
One possibility would be to consider the Karush-Kuhn-Tucker conditions by managing active constraints and integrating a piecewise smooth system of differential equations. However, due to the fact that in our experiments this situation has never occurred, we expect to observe it only exceptionally. For this reason we treat this situation by simply adding a suitable penalty term to the functional $F_{\varepsilon}(E)$,
$F_{\varepsilon}^{c}(E)=F_{\varepsilon}(E)+\frac{c}{\varepsilon}\,P_{\varepsilon}(E)$
for increasing $c\gg\varepsilon$, with
(21) $P_{\varepsilon}(E)=\frac{1}{2}\sum\limits_{(i,j)\in\mathcal{E}}\left(a_{ij}+\varepsilon e_{ij}-\delta\right)_{-}^{2}$
where $(p)_{-}=\min\left(p,0\right)$. Differentiating $P_{\varepsilon}$ yields
$\nabla P_{\varepsilon}(E)=\varepsilon\sum\limits_{(i,j)\in\mathcal{E},a_{ij}+\varepsilon e_{ij}<\delta}\left(a_{ij}+\varepsilon e_{ij}-\delta\right)_{-}$
which has to be included in the gradient system (19). The modified matrix-valued ODE is
(22) $\dot{E}=-G_{\varepsilon}^{c}(E)+\gamma E\quad\hbox{ with }\quad\gamma=\langle E,G_{\varepsilon}^{c}(E)\rangle$
where
(23) $G_{\varepsilon}^{c}(E)=G_{\varepsilon}(E)+\frac{c}{\varepsilon}\,\nabla P_{\varepsilon}(E).$
Whenever a trajectory $E(t)$ determines a negative entry of $A+\varepsilon E(t)$, the gradient system tends to recover non-negativity, with an effect which is proportional to $c$.
Similarly to Theorems 5.1 and 5.2 we have the following results, whose proofs are omitted since they are analogous to those of previous theorems.
Theorem 5.3 (Monotonicity of the penalized functional).
Let $E(t)$ of unit Frobenius norm satisfy the differential equation (22) with arbitrary but fixed $c>0$ and $G^{c}_{\varepsilon}(E)$ given by (23). Then, the functional $F_{\varepsilon}^{c}(E(t))$ decreases monotonically with respect to $t$: $\dot{F}^{c}_{\varepsilon}(E(t))\leq 0.$
Theorem 5.4 (Stationary points of the penalized functionl).
The following statements are equivalent along solutions of (22):
  1. (1)
    ${\dot{F}}_{\varepsilon}^{c}=0$.
  2. (2)
    $\dot{E}=0$.
  3. (3)
    $E$ is a real multiple of $G_{\varepsilon}^{c}(E)$.
5.6. Undirected graphs

### v2

@@ 5.2. Structured gradient of the functional (7)
with an appropriate matrix-valued function $Z_{\varepsilon}$, which – in a natural way – is interpreted as the constrained gradient of $F_{\varepsilon}$ (with respect to the Frobenius inner product). Projecting it onto the set of real matrices with sparsity pattern induced by ${\mathcal{E}}$ (first constraint) and unit norm (second constraint), we obtain the sought fully constrained gradient.
To address the undirected case, we define for $A\in\mathbb{R}^{n\times n}$,
$\mathrm{Sym}(A)\big{|}_{ij}:=\frac{a_{ij}+a_{ji}}{2}$
@@
Lemma 5.1 (Structured gradient of the functional  –  directed graphs).
@@
(15) $\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=\eta\Big{\langle}G_{\varepsilon}(E),\dot{E}\Big{\rangle},$
where $\eta>0$ is a suitable scaling factor, with the structured gradient $G_{\varepsilon}(E)$, having the sparsity pattern determined by the set of edges $\mathcal{E}$)
(16) $G_{\varepsilon}(E)=P_{\mathcal{E}}\left(R\right)$
where $R$ is the free gradient:
$R$ $=\sum\limits_{k=1}^{m}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)\Bigl{(}\left(\mathrm{v}_{i_{k}}-\langle\mathrm{v}\rangle_{m}\right)(A_{\varepsilon}^{\#})^{\top}\mathrm{v}\mathrm{v}^{\top}-(A_{\varepsilon}^{\#})^{\top}\left({\rm u}_{i_{k}}-\bar{\rm u}\right)\mathrm{v}^{\top}\Bigr{)}.$
@@
Theorem 4.3 applied to $A+\varepsilon E(t)$ yields
@@
This gives the structure-constrained gradient $G_{\varepsilon}(E)=P_{\mathcal{E}}(R)$ (see (16)). ∎
Lemma 5.2 (Fully constrained gradient of the functional – undirected graphs).
Under the same assumptions of Lemma 5.1 to which is added the hypothesis that the smooth path $E(t)$ is also symmetric, we have that the fully constrained gradient is still given by (15) with
Proof.
The proof is similar to that of Lemma 5.1. Here $A_{\varepsilon}^{\#}=A_{\varepsilon}^{\dagger}$; then, by the well-known property $\mathrm{v}^{\top}A_{\varepsilon}^{\dagger}=0$, we first get
$R=-(A_{\varepsilon}^{\#})^{\top}\bigl{(}{\rm u}_{i_{k}}-\bar{\rm u}\bigr{)}\mathrm{v}^{\top}.$
Next, and including the symmetry constraint (ic) and thus the symmetry of $\dot{E}$, we get
$\big{\langle}R,\dot{E}\big{\rangle}=\big{\langle}P_{\mathcal{E}}(R),\dot{E}\big{\rangle}=\big{\langle}\mathrm{Sym}\left(P_{\mathcal{E}}(R)\right)+\mathrm{Skew}\left(P_{\mathcal{E}}(R)\right),\dot{E}\big{\rangle}=\big{\langle}\mathrm{Sym}\left(P_{\mathcal{E}}(R)\right),\dot{E}\big{\rangle}$
where for a matrix $B$, $\mathrm{Skew}(B)={(B-B^{\top})}/{2}$. Orthogonality of symmetric and skew-symmetric matrices proves the result, i.e. gradient (16) symplifies to (18). ∎
5.3. Constrained gradient of the functional
In this section we take into account the norm constraint on $E$ and the inequality constraint on $A+\varepsilon E$. One possibility would be to use a penalization method; this, however, would have the disadvantage that the potential violation of the inequality constraints might have an impact on the Perron eigenvector and on its continuity. For this reason we opt for a Lagrangian approach.
Admissible directions
We recall the considered constraints on $E$:
  • (a)
    $E$ is of unit Frobenius norm.
  • (b)
    $E=P_{\mathcal{E}}(E)$.
  • (c)
    $(A+\varepsilon E)_{ij}\geq\delta$ for $(i,j)\in\mathcal{E}$.
  • (d)
    $E$ is symmetric (only for undirected graphs).
Since $E(t)$ is of unit Frobenius norm by condition (a), we have
$0=\frac{1}{2}\,\frac{d}{dt}\lVert E(t)\rVert^{2}=\langle E(t),\dot{E}(t)\rangle.$
To fulfill the non-negativity condition (c) of the matrix $A+\varepsilon E(t)$, we need that
$\dot{e}_{ij}(t)\geq 0\quad\mbox{for all}\ (i,j)\ \mbox{with}\ a_{ij}+\varepsilon e_{ij}(t)=\delta\geq 0.$
Conditions (b) and (d) are satisfied if the same condition holds for $\dot{E}$. These four conditions are in fact also sufficient for a matrix to be the time derivative of a path of $\varepsilon$-feasible matrices.
Hence, for every $\varepsilon$-feasible matrix $E$, a matrix $Z=(z_{ij})\in\mathbb{R}^{n\times n}$ is the derivative of some path of $\varepsilon$-feasible matrices passing through $E$ if and only if the following four conditions are satisfied:
  • (a’)
    $\langle E,Z\rangle=0$.
  • (b’)
    $Z=P_{\mathcal{E}}(Z)$.
  • (c’)
    $P_{{\mathcal{E}}_{0}}(Z)\geq 0$.
  • (d’)
    $Z$ is symmetric (only for undirected graphs) .
Here $\mathcal{E}_{0}={\mathcal{E}}_{0}(\varepsilon E)$ is the set
$\mathcal{E}_{0}:=\{(i,j)\in\mathcal{E}:\,a_{ij}+\varepsilon e_{ij}=\delta\}.$
Condition (c’) implies that $z_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}$.
Admissible direction of steepest descent
To determine the admissible direction $\dot{E}$ of steepest descent from $E$, we therefore consider the following optimization problem for $G=G_{\varepsilon}(E)$:
(19) $\min_{Z}\langle G,Z\rangle\quad\text{subject to}$
(20) $\begin{cases}\langle E,Z\rangle=0\,,\\ z_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,\\ z_{ij}=z_{ji}\text{ for all }(i,j)\quad\text{(for undirected graphs})\,,\\ \langle Z,Z\rangle=1\,.\end{cases}$
The additional constraint $\lVert Z\rVert=1$ just normalizes the descent direction. Problem (19) has a quadratic constraint.
We now formulate a quadratic optimization problem with linear constraints, which is equivalent to (19).
An equivalent quadratic optimization problem with linear constraints
(21) $\min_{Z}\langle Z,Z\rangle\quad\text{subject to}$
(22) $\begin{cases}\langle E,\dot{E}\rangle=0\,,\\ \dot{E}_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,\\ z_{ij}=z_{ji}\text{ for all }(i,j)\quad\text{(for undirected graphs})\,,\\ \langle G,Z\rangle=-1\,.\end{cases}$
Problems (19) and (21) are equivalent in the sense that Problem (21) yields the same solution, i.e. descent direction, provided that a strict descent direction exists, i.e., satisfying $\langle G,Z\rangle<0$ and the constraints (a’)–(d’). The equivalence is a consequence of the fact that when $\langle G,Z\rangle<0$, there exists a scaling factor $\alpha>0$ s uch that $\langle G,\alpha Z\rangle=-1$.
Apart from the normalization, problems (19) and (21) yield the same KKT (Karush–Kuhn–Tucker) conditions. Since the objective function $\langle Z,Z\rangle$ of problem (21) is convex and all constraints are linear, the KKT conditions are not only necessary but also sufficient conditions ([10, Theorem 9.4.1]), that is, a KKT point is already a solution of the optimization problem.
The solution of (21) satisfies the KKT conditions (see, e.g., [11, Ch. 2.2])
$\kappa Z=-G-\gamma E+\sum_{(i,j)\in\mathcal{E}_{0}}\mu_{ij}{\rm u}_{i}{\rm u}_{j}^{T}\,,$
$Z_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,$
$\langle E,Z\rangle=0\,,$
$\langle Z,Z\rangle=1\,,$
[…]
$\mu_{ij}z_{ij}=0\text{ for all }(i,j)\in\mathcal{E}_{0}\,,$
$\mu_{ij}\geq 0\text{ for all }(i,j)\in\mathcal{E}_{0}\,.$
where ${\rm u}_{\ell}$ is the $\ell$-th unit versor, $\gamma$ is determined by the normalization $\langle E,Z\rangle=0$ and $\kappa>0$ is determined by the normalization $\lVert Z\rVert^{2}=\langle Z,Z\rangle=1$.
5.4. Constrained gradient flow
The gradient flow of $F_{\varepsilon}$ under the constraints (a)–(d) is given by
(24) $\dot{E}(t)=Z(t),$
where $Z(t)$ solves the KKT system (23) with $G=G_{\varepsilon}(E(t))$ under the constraints (a’)–(d’). Note that the set of edges ${\mathcal{E}}_{0}(t)={\mathcal{E}}_{0}(\varepsilon E(t))$.
Lemma 5.3.
On an interval where ${\mathcal{E}}_{0}(t)$ does not change, the gradient system becomes, with $P^{+}=P_{{\mathcal{E}}\setminus{\mathcal{E}}_{0}}$,
@@ 5.4. Constrained gradient flow
(25) $\dot{E}=-P^{+}G_{\varepsilon}(E)+\gamma P^{+}E\quad\hbox{ with }\quad\gamma=\frac{\langle G_{\varepsilon}(E),P^{+}E\rangle}{\|P^{+}E\|^{2}}.$
Proof.
The positive Lagrange multipliers $\mu_{ij}>0$ just have the role to ensure that $\dot{e}_{ij}=0$. With $G=G_{\varepsilon}(E)$, the gradient system therefore reads
(26) $\dot{E}=P^{+}(-G+\gamma E),$
where $\gamma$ is determined from the constraint $\langle E,\dot{E}\rangle=0$. We then have
$0=\langle E,\dot{E}\rangle=\langle E,P^{+}(-G+\gamma E)\rangle=-\langle P^{+}E,G\rangle+\gamma\langle P^{+}E,P^{+}E\rangle,$
and the result follows. ∎
@@ 5.5. Monotonicity and stationary points
The following result shows the monotonicity of the functional $F_{\varepsilon}(E(t))$ with respect to $t$ and provides a characterization of the stationary points. We observe here that monotonicity follows directly from the construction of the gradient system (25) (for the time being, we assume that the non-negativity constraint does not activate so that penalization is not necessary, which is what we observe in all the tests we have done).
@@
Furthermore, the following statements are equivalent along solutions of (25):
@@
    $P^{+}E$ is a real multiple of $P^{+}G_{\varepsilon}(E)$.
@@
We prove the result in three steps.
  • (1) implies (3): Using Lemma 5.1 and (25) we obtain, with $G=G_{\varepsilon}(E)$,
    (27) $\frac{1}{\varepsilon}\,\dot{F}_{\varepsilon}=\langle G,\dot{E}\rangle=\langle G,-P^{+}G+\gamma P^{+}E\rangle=-\|P^{+}G\|^{2}+\frac{\langle P^{+}G,P^{+}E\rangle^{2}}{\|P^{+}E\|^{2}}.$
    By means of the strong form of the Cauchy–Schwarz inequality we obtain monotonicity. Furthermore we have that
    $\frac{d}{dt}F_{\varepsilon}\left(E(t)\right)=0\qquad\Longrightarrow\qquad\mbox{$P^{+}E$ is a multiple of $P^{+}G$}.$
  • (3) implies (2): This is a direct consequence of the vanishing of the r.h.s. of (25).
  • (2) implies (1): This is also a direct consequence of Lemma 5.1.
@@ 5.6. Underlying low-rank property of extremizers
For any graph, we have a number of ODEs which is given by the nonzero entries in the adjacency matrix associated to the graph. Hence, for a dense graph we may have up to $n^{2}$ ODEs. If inequality constraints are inactive, the gradient system (25) is such that $P^{+}$ is the identity and the gradient system therefore reads
(28) $\dot{E}=-G+\gamma E,\qquad\gamma=\langle G,E\rangle.$
where $G=G_{\varepsilon}(E)$. In this case extremizers (stationary points) are characterized by an underlying low-rank property, being $G$ the projection of a low-rank matrix onto the sparsity-pattern of the weighted adjacency matrix. Clearly, if $\varepsilon<\min_{ij}a_{ij}$, inequality constraints cannot be active so that we can ignore them and solve the following system of ODEs (28). Moreover, in all our experiments we have never observed activation of inequality constraints associated to nonnegativity.
@@
Following the lines of [17], we are able to exploit the underlying low-rank structure of the solutions of the optimization problem we consider. By Theorem 5.1 we have that the stationary points which we aim to compute are proportional to the gradient $G$, which - as we have seen - is the projection of a rank-$1$ matrix. Proceeding similarly to [17], by writing $E=P_{\mathcal{E}}(H)$ for a rank-$1$ matrix $H$, we are able to obtain a new system of ODEs for $H$, whose dynamics is constrained to the manifold of rank-$1$ matrices. We can prove that the stationary points of the two systems of ODEs are in a one-to-one correspondence and also that the system of ODEs for $H$, even if it loses the gradient system structure, converges - at least locally - to the stationary points. We omit a full derivation for sake of conciseness and refer the reader to [17] for details.
5.7. Including a linear constraint on the entries of the adjacency matrix
Suppose now that a few weights associated with the edge set $\mathcal{E}^{\prime}\subseteq\mathcal{E}$ are related by a linear relationship, that is
(29) $\sum_{(i,j)\in\mathcal{E}}b_{ij}a_{ij}=c,$
with $c,b_{ij}\in\mathbb{R}$ for all $(i,j)\in\mathcal{E}^{\prime}$ and $b_{ij}=0$ for all $(i,j)\in\mathcal{E}\setminus\mathcal{E}^{\prime}$. Then for the perturbed matrix $A+\varepsilon E(t)$ we have to impose for all $t$
$\sum_{(i,j)\in\mathcal{E}}b_{ij}e_{ij}(t)=0\quad\Longrightarrow\quad\sum_{(i,j)\in\mathcal{E}}b_{ij}\dot{e}_{ij}(t)=0,$
that is, we have to impose the scalar product $\langle b,\dot{e}(t)\rangle=0$ for the associated vectors $b=\left(b_{ij}\right)$ and $e=\left(e_{ij}\right)={\rm vec}(E)$, which represents a column vector formed by the nonzero entries of $E$ (which are defined by the sparsity pattern).
This implies that we have to modify the r.h.s. in the matrix ODE (24) in the inner iteration, which we write in short as $\dot{E}=Z$. Writing it in vectorized form as $\dot{e}=z$, with $z={\rm vec}(Z)$, we modify it as
(30) $\dot{e}=z-\frac{1}{\|b\|^{2}}\langle z,b\rangle b,$
that is, we orthogonally project the vector field $z$ onto the linear space ${\rm Ker}(b)=\langle b,\cdot\rangle=0$. Indeed, this implies
$\langle\dot{e},b\rangle=\langle z,b\rangle-\frac{1}{\|b\|^{2}}\langle z,b\rangle\langle b,b\rangle=0,$
as desired; it follows that the property (29) is conserved along the flow of the modified ODE. Equation (24) is in vector form, but it can be rewritten directly in matrix form.
6. Undirected graphs